OpenStudy (anonymous):

Solve the following equation by transforming it into a perfect square trinomial. 3x2 + 18x = 21 A. {–19, 13} B. {–25, 7} C. {–7, 1} D. {–13, –5}

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

An equation in which one side is a perfect square trinomial can be easily solved by taking the square root of each side.

OpenStudy (anonymous):

What

OpenStudy (anonymous):

Okay lets work this out ..

OpenStudy (anonymous):

1st Keep all x related terms on one side and move the constant to the right.

OpenStudy (anonymous):

next Balance the equation.

OpenStudy (anonymous):

I'm lost

OpenStudy (anonymous):

Ok so consider the equation: 3x^2 + 18x = 21 Our solution strategy will be to get a perfect square trinomial of the form a^2 + 2ab + b^2 = (a + b)^2

OpenStudy (anonymous):

on the left hand side of the equation.